### v1

# Brought a Gun to a Knife fight: Modern VFM Baselines Outgun Specialized Detectors on In-the-Wild AI Image Detection

Abstract
While specialized detectors for AI-generated images excel on curated benchmarks, they fail catastrophically in real-world scenarios, as evidenced by their critically high false-negative rates on ‘in-the-wild’ benchmarks. Instead of crafting another specialized ‘knife’ for this problem, we bring a ‘gun’ to the fight: a simple linear classifier on a modern Vision Foundation Model (VFM). Trained on identical data, this baseline decisively ‘outguns’ bespoke detectors, boosting in-the-wild accuracy by a striking margin of over 20%.
Our analysis pinpoints the source of the VFM’s ‘firepower’: First, by probing text-image similarities, we find that recent VLMs (e.g., Perception Encoder, Meta CLIP2) have learned to align synthetic images with forgery-related concepts (e.g., ‘AI-generated’), unlike previous versions. Second, we speculate that this is due to data exposure, as both this alignment and overall accuracy plummet on a novel dataset scraped after the VFM’s pre-training cut-off date, ensuring it was unseen during pre-training. Our findings yield two critical conclusions: 1) For the real-world ‘gunfight’ of AI-generated image detection, the raw ‘firepower’ of an updated VFM is far more effective than the ‘craftsmanship’ of a static detector. 2) True generalization evaluation requires test data to be independent of the model’s entire training history, including pre-training.
Index Terms—  AIGI Detection, Multimedia Forensics
1 Introduction
The field of Artificial Intelligence Generated Images (AIGI) has seen explosive growth in recent years, with a particular focus on the creation of highly realistic synthetic images through advanced generative models. These synthetic images have significantly fueled the proliferation of misinformation, creating serious threats to both societal security and individual privacy.
As a result, a central challenge for AIGI detection is building models with a strong generalization ability, allowing them to effectively identify and verify images generated by a variety of unknown methods. To address this, researchers have developed various forensics-specialized detectors, such as incorporating additional augmented samples (e.g., full or partial image reconstructions) [1, 2, 3] or designing improved training strategies [4]. These approaches have already pushed the performance on recent AIGI benchmarks to impressive levels. Consequently, the research focus has begun to shift toward the more challenging in-the-wild datasets. However, such real-world scenarios remain highly challenging, as most forensic methods fail to reach satisfactory performance, particularly on the Chameleon [5], as illustrated in Fig. 1.
Fig. 1: Comparison of detection performance between modern VFMs and state-of-the-art forensics-specialized detectors on the GenImage and Chameleon.
In this work, we show that a surprisingly simple baseline: a linear classifier using a modern VFM as a feature extractor, training the same examples without any augmentation, already outperformed even the most advanced forensic detectors on in-the-wild datasets, improving accuracy by over 20%. In particular, a new generation of foundation models, spanning both Vision-Language (e.g., PE[6], Meta CLIP-2[7]) and self-supervised architectures (e.g., DINOv3[8]), though not being specifically designed for forensics, exhibit strikingly superior performance compared to both earlier VFMs and forensics-specific detectors. This raises a critical question: does their success arise from genuinely stronger representations, or from inadvertent exposure to synthetic data during pre-training?
To investigate this issue, we conduct two analyses. First, focusing on the VLM-based backbones, we probe text-image similarities and find they have already aligned synthetic images with forgery-related concepts (e.g., ‘AI-generated’) well, a notable characteristic absent in earlier models. This provides a clear concept clue for their success. Second, to verify the broader hypothesis of data exposure for all tested VFMs, we collected a novel dataset of images generated after their pre-training cut-off date. As predicted, nearly all top models exhibit a noticeable performance drop. Thus, their remarkable performance appears to be a direct consequence of the model’s inadvertent training on the concept of contemporary AIGI within its pre-training data. This effectively equips them with a ’built-in’ effectiveness to in-the-wild content, allowing these general-purpose models to decisively outgun specialist detectors explicitly designed for this forensic task.
In summary, our main contributions are:
  • We establish the superiority of simple, modern VFM baselines over specialized detectors for in-the-wild scenarios. We argue that for real-world effectiveness, the best strategy is to leverage the most up-to-date pre-trained knowledge, even if it stems from data exposure.
  • By identifying data exposure as the primary source of this success, we reveal a fundamental flaw in static benchmarks and thus advocate for a crucial shift towards dynamic, continuously updated evaluation protocols. This will ensure they remain verifiably unseen, providing a more accurate and future-proof measure of a model’s ability to handle genuinely novel threats.
2 Evaluation
2.1 Experimental Setup
Models. We evaluate a diverse set of Vision Foundation Models (VFMs), categorized into three groups: 1) Modern VFMs released after early 2025 (Meta CLIP-2 [7], Pe [6], SigLIP-2 [9]); 2) their direct Predecessors (CLIP [10], Meta CLIP [11], SigLIP [12]); and 3) Self-supervised Models (DINOv3 [8] and DINOv2 [13]). For all VFMs, we use their largest, officially released weights.
Evaluation Datasets. Our evaluation focuses on established in-the-wild datasets, each designed to capture different facets of real-world scenarios. We primarily use four benchmarks which can be categorized by their collection sources:
  • Social Media Sourced: Datasets like WildRF [14] and SocialRF [15] are collected from broad social media platforms such as X (formerly Twitter), Facebook, and Reddit. Their collection process typically involves querying common hashtags associated with authentic content (e.g., ‘#photography‘, ‘#nofilter‘) and AI-generated content (e.g., ‘#AIart‘, ‘#midjourney‘).
  • AI Art Community Sourced: In contrast, datasets like Chameleon [5] and CommunityAI [15] are curated from specialized AI art communities like ArtStation and Civitai. These benchmarks prioritize high-realism, high-quality images from diverse models, often reflecting the cutting edge of generative art.
Together, these datasets provide a comprehensive and challenging testbed for evaluating a model’s true generalization capabilities.
Training and Evaluation Protocol. For each VFM, we train a single linear classifier on its frozen ‘[CLS]’ token features, using the GenImage [16] (SD v1.4 subset) without any augmentations. The resulting classifiers are then evaluated on all in-the-wild datasets. For comparison against SOTA forensic detectors, we report their performance directly from the original publications, as they often rely on specialized training strategies incompatible with our unified setup.
Table 1: Comparison results on GenImage and Chameleon.
Method GenImage Chameleon Avg. R.Acc F.Acc Avg. CNNSpot (CVPR’20) [17] 64.2 99.7 0.08 56.9 FreqNet (AAAI’24) [18] 86.8 87.1 13.7 55.6 GramNet (CVPR’20) [19] 69.9 99.6 4.76 58.9 UnivFD (CVPR’23) [1] 88.8 97.8 3.18 57.2 NPR (ECCV’24) [20] 88.6 98.7 2.20 57.3 AIDE (ICLR’25) [5] 86.8 98.5 5.04 58.4 PPL (arxiv’25) [21] 97.2 68.1 64.7 66.6 OMAT (arxiv’25) [22] 94.6 90.2 33.9 66.1 NPLB (arxiv’25) [4] 97.1 79.4 65.5 73.4 CLIP (2021.2.26) 84.3 70.7 35.5 55.6 Siglip (2023.3.27) 88.6 70.8 49.3 61.6 Siglip2 (2025.2.21) 92.0 88.9 77.0 83.8 Meta CLIP (2023.9.28) 76.6 37.6 91.5 60.7 Meta CLIP-2 (2025.7.29) 89.5 84.8 87.9 91.8 PE (2025.4.17) 93.8 97.0 94.8 96.1 DINOv2 (2023.10.26) 85.3 62.9 58.1 60.8 DINOv3 (2025.8.14) 96.3 93.7 90.5 92.4
2.2 Performance on Curated Benchmarks vs. Chameleon
We first evaluate a wide range of specialized forensic detectors and our VFM baselines on both the curated GenImage benchmark and the challenging in-the-wild Chameleon dataset. The results, presented in Table 1, reveal a stark performance dichotomy.
On the GenImage benchmark, many recent specialized detectors like PPL (97.2%) and NPLB (97.1%) achieve impressive accuracy, often leveraging complex training strategies. However, this success is brittle and fails to translate to the more realistic Chameleon dataset. A dramatic performance collapse is observed across most specialized methods; traditional detectors like CNNSpot and UnivFD fail entirely, with Fake Accuracy (F.Acc) dropping to near zero. This reveals a critical disconnect between benchmark performance and the generalization required for real-world application.
In stark contrast, modern VFM baselines demonstrate exceptional effectiveness. Despite their simple training, models like PE, Meta CLIP-2, and DINOv3 surpass 90% accuracy on Chameleon, decisively outperforming the entire class of specialized detectors. This suggests that for practical, in-the-wild scenarios, the general-purpose representations of modern VFMs are fundamentally superior to task-specific designs.
2.3 Validation on Additional In-the-Wild Datasets
To further validate these findings, we selected several additional representative in-the-wild datasets (i.e., WildRF, CommunityAI and SocialRF) and conducted the same comparative experiments, as shown in Tab. 2 and Tab. 3. Nearly all forensic detectors lost their detection capabilities on WildRF, with only a few exceptions achieving good performance. However, all vision encoders surpassed a 70% accuracy rate, and DINOv3 even reached an impressive 96.4%. Similarly, Tab. 3 shows that all forensic-specialized detectors failed to perform on these datasets, while most vision encoders maintained excellent detection capabilities. For instance, Perception achieved a 96% accuracy on the CommunityAI dataset, and DINOv3 reached 93% on SocialRF.
Table 2: Comparison results on WildRF.
Method WildRF Facebook Reddit Twitter Avg. NPR (ECCV’24) [20] 78.1 61.0 51.3 63.5 UnivFD (CVPR’23) [1] 49.1 60.2 56.5 55.3 FatFormer (CVPR’24) [23] 54.1 68.1 54.4 58.9 SAFE (KDD’25) [24] 50.9 74.1 37.5 57.2 C2P-CLIP (AAAI’25) [25] 54.4 68.4 55.9 59.6 AIDE (ICLR’25) [5] 57.8 71.5 45.8 58.4 DRCT (ICML’24) [2] 46.6 53.1 55.2 50.6 AlignedForensics (ICLR’25)[26] 89.4 69.1 81.8 80.1 DDA (arxiv’25) [27] 94.4 94.3 96.8 95.1 CLIP (2021.2.26) 77.2 72.3 68.9 72.8 Siglip (2023.3.27) 58.4 66.5 72.2 65.7 Siglip2 (2025.2.21) 79.9 84.7 89.6 84.7 Meta CLIP (2023.9.28) 63.8 74.1 77.9 71.9 Meta CLIP-2 (2025.7.29) 70.9 76.7 81.2 76.3 Pe (2025.4.17) 81.6 84.9 88.8 85.1 DINOv2 (2023.10.26) 71.3 71.0 71.3 71.2 DINOv3 (2025.8.14) 94.7 96.7 97.9 96.4
Table 3: Comparison results on SocialRF and CommunityAI.
Method SocialRF CommunityAI R.Acc F.Acc Avg. R.Acc F.Acc Avg. CNNspot (CVPR’20) [17] 94.7 7.5 51.1 97.3 5.3 51.3 GramNet (CVPR’20) [19] 92.6 11.5 52.1 99.0 6.2 52.6 UnivFD (CVPR’23) [1] 53.3 55.5 54.4 82.8 51.2 67.0 FreqNet (AAAI’24) [18] 68.5 39.3 53.9 98.9 12.2 53.6 NPR (ECCV’24) [20] 96.3 21.9 59.1 99.9 8.2 54.05 AIDE (ICLR’25) [5] 97.2 18.4 57.8 99.0 9.3 54.15 SAFE (KDD’25) [24] 99.6 16.4 58.0 100.0 8.5 54.25 CLIP (2021.2.26) 41.0 77.3 59.1 69.8 42.4 56.1 Siglip (2023.3.27) 47.3 62.3 54.8 49.7 55.2 52.4 Siglip2 (2025.2.21) 64.8 90.2 77.5 83.1 86.3 84.7 Meta CLIP (2023.9.28) 40.5 86.4 63.5 35.5 93.4 64.5 Meta CLIP-2 (2025.7.29) 41.6 92.4 67.0 91.7 94.9 93.3 PE (2025.4.17) 52.1 98.1 75.1 96.6 97.5 97.1 DINOv2 (2023.10.26) 60.3 69.5 64.9 60.6 56.2 58.4 DINOv3 (2025.8.14) 88.7 98.0 93.3 95.6 95.0 95.3
3 Analysis
3.1 Data Exposure Hypothesis: A Test on Post-Cutoff Data
Our primary speculation is that the superior in-the-wild performance of modern VFMs is a direct result of exposure to contemporary synthetic media during pre-training, rather than an intrinsic leap in generalization. To rigorously test this, we constructed a new dataset, WebAIG-25, designed to serve as a decisive blind test. We built it by collecting synthetic samples posted on the subreddit ‘r/midjourney‘ and real-world photographs from Unsplash. Crucially, all samples were uploaded by users between August 1, 2025, and September 10, 2025. This specific timeframe places the data after the pre-training data cutoff for most models under evaluation, making WebAIG-25 a powerful tool to disentangle true generalization from learned artifacts.
Table 4: Comparison results on WebAIG-25.
Method WebAIG-25 R.Acc F.Acc Avg. CLIP (2021.2.26) 52.1 71.1 61.6 Siglip (2023.3.27) 46.7 44.1 45.4 Siglip2 (2025.2.21) 84.8 55.8 70.3 Meta CLIP (2023.9.28) 59.7 55.6 57.7 Meta CLIP-2 (2025.7.29) 80.1 75.4 77.7 PE (2025.4.17) 100.0 91.7 95.9 DINOv2 (2023.10.26) 57.0 54.2 55.6 DINOv3 (2025.8.14) 98.6 91.4 95.0
The results on this ‘post-cutoff’ dataset, presented in Table 4, are revealing and offer strong support for our hypothesis. The most telling evidence comes from the modern VFM families. While the modern MetaCLIP-2 (77.7%) still outperforms its predecessor MetaCLIP (57.7%), its accuracy represents a stark drop from the 91.8% it achieved on Chameleon. A similar, significant drop is also observed for Siglip2. This pattern strongly suggests their high performance on older ‘in-the-wild’ sets is less an indicator of true generalization and more a result of prior data exposure. It is this very performance drop on verifiably unseen data that serves as the most compelling evidence, confirming that data exposure—not a fundamental leap in abstract generalization—is the dominant factor driving the remarkable success of modern VFMs on established benchmarks.
Table 5: Comparison of Text–Image Similarities on In-the-Wild Dataset
Method Chameleon SocialRF Top-1 Top-2 Top-1 Top-2 Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score CLIP (2021.2.26) modern design 0.628 portrait 0.345 forged 0.332 urban 0.253 Siglip (2023.3.27) unaltered 0.318 original 0.212 edited 0.232 original 0.222 Siglip2 (2025.2.21) genuine 0.385 urban 0.250 portrait 0.202 vintage Meta CLIP (2023.9.28) AI generated 0.678 deepfake 0.091 AI generated 0.902 original 0.024 Meta CLIP-2 (2025.7.29) AI generated 0.828 deepfake 0.064 AI generated 0.924 deepfake 0.038 PE (2025.4.17) AI generated 0.861 deepfake 0.021 AI generated 0.943 deepfake 0.031 Method CommunityAI WebAIG-25 Top-1 Top-2 Top-1 Top-2 Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score CLIP (2021.2.26) portrait 0.346 nature 0.326 urban 0.332 midjourney_images 0.260 Siglip (2023.3.27) real 0.281 AIGIBench 0.283 fake 0.278 original 0.246 Siglip2 (2025.2.21) urban 0.209 portrait 0.201 urban 0.212 portrait 0.194 Meta CLIP (2023.9.28) AI generated 0.726 deepfake 0.086 midjourney_images 0.604 AI generated 0.284 Meta CLIP-2 (2025.7.29) AI generated 0.858 deepfake 0.041 midjourney_images 0.621 AI generated 0.308 PE (2025.4.17) AI generated 0.878 CommunityAI 0.029 midjourney_images 0.722 AI generated 0.218
3.2 Probing VFM’s Latent Alignment with Forgery Concepts
To investigate the source of the superior in-the-wild performance observed in modern foundation models, we speculate that recent Vision-Language Models (VLMs) may have learned to associate synthetic images with forgery-related semantic concepts during pre-training. This would provide a powerful, high-level cue for detection that earlier models and non-language-aligned models like DINOv3 might lack.
To test this hypothesis, we designed a text-image similarity probing experiment exclusively for the VLM-based backbones. Specifically, we randomly sample images from several in-the-wild datasets. For each image, we compute its visual embedding using the VLM’s image encoder and measure its cosine similarity to a set of textual embeddings generated by the corresponding text encoder. These textual concepts are categorized into three groups:
  • Real and Fake-Related Keywords: [‘fake’, ‘real’, ‘AI generated’, ‘authentic’, ‘manipulated’, ‘synthetic’, etc],
  • Redundant Keywords: [‘sunset’, ‘landscape’, ‘portrait’, ‘abstract art’, ‘technology’, ‘nature’, etc],
  • Training Dataset Names: [‘GenImage’, ‘ADM’, ‘BigGAN’, ‘glide’, ‘Midjourney’, etc].
We randomly sampled 2,000 images from each in-the-wild dataset. Then calculated the similarity between these images and our pre-defined text pool, outputting the most textually similar terms for each image.
Table 5 reveals a significant generational divide in semantic understanding. Older models like CLIP and Siglip consistently fail to associate fake images with forgery-related concepts, instead matching them to generic terms like ‘portrait’ or ‘urban’. Their latent space essentially lacks a concept of "fakeness."
In stark contrast, modern VLMs like MetaCLIP-2 and Pe demonstrate a remarkably strong and consistent alignment. Across both established in-the-wild datasets and our new WebAIG-25, their top-matched concepts for fake images are overwhelmingly forgery-related terms like "AI generated" or "midjourney_images". This powerful semantic alignment, learned through exposure to contemporary web data, serves as a decisive shortcut for detection. It confirms that modern VLMs excel not just by recognizing visual artifacts, but by understanding the very concept of a "fake" image.
3.3 Visualization
The t-SNE visualizations in Fig. 2 offer a compelling qualitative explanation for the quantitative performance gaps in Tab. 1. On the curated GenImage dataset, both Meta CLIP-2 and CLIP exhibit similarly entangled feature spaces, which aligns with the comparable, high performance of both our VFM baselines (89.5%) and CLIP-based detectors like UnivFD (88.8%). However, this parity vanishes on the in-the-wild Chameleon dataset. There, the performance of UnivFD collapses to 57.2%, a failure visually mirrored in its CLIP backbone’s chaotic and inseparable feature space. In stark contrast, MetaCLIP2’s near-perfect accuracy (91.8%) is directly explained by its t-SNE plot, which reveals a remarkably clean separation between real and fake clusters. This direct link between feature separability and downstream performance confirms that MetaCLIP2’s advantage stems from a fundamentally more discriminative representation for in-the-wild content, acquired through pre-training exposure (Fig. 2).
Fig. 2: T-SNE Visualization of CLIP and Meta CLIP-2 on GenImage and Chameleon.
4 Conclusion
This paper demonstrated that for in-the-wild AI-generated image detection, a simple linear classifier built upon a modern Vision Foundation Model (VFM) decisively outperforms specialized forensic detectors, akin to bringing a gun to a knife fight. We attribute this stark performance gap not to architectural novelty but to the VFM’s pre-training on vast, contemporary web data that includes synthetic media. Our findings have two critical implications: first, for practical forensics, leveraging the evolving representations of the latest foundation models is a more robust strategy than designing static detectors. Second, we call for a stricter evaluation protocol where test data is guaranteed to be novel to a model’s entire pre-training corpus to truly assess generalization.